OpenStudy (anonymous):

\[y''+3y=0\] \[y(0)=1\] \[y'(0)=3\] \[y(0)=c_1cos\sqrt3(0)-c_2sin\sqrt3(0)=1\]

OpenStudy (anonymous):

Do you know what \(\cos(0)\) and \(\sin(0)\) are?

OpenStudy (anonymous):

yes, cos(0)=1

OpenStudy (anonymous):

why not just find the derivative of y(x)?

OpenStudy (anonymous):

\[c_2\sin\sqrt({3}*0)=0\] leaves \[c_1\cos(\sqrt{3}*0)=1\] \[c_1*1=1\] \[c_1=1\] \[y(x)=c_1\cos(x\sqrt{3})-c_2\sin(x\sqrt{3})\] \[y'(x)=-c_1\sqrt{3}sin(x\sqrt{3})-c_2\sqrt{3}cos(x\sqrt{3})\] i hope i didnt derive it incorrectly anyways \[f'(0)=3=-c_1\sqrt{3}sin(0\sqrt{3})-c_2\sqrt{3}cos(0\sqrt{3})\] substitute in c_1 and solve for c_2

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

so would \[c_2=\frac{-3}{\sqrt3}\]